One way in whith the useful metal copper is produced is by dissolving the mineral azurite, which contains copper(Il) carbonate, in concentrated sulfuric acid. The sulfuric acd reacts with the copper(II) carbonate to produce a blue solution of copper(II) sulfate. Scrap iron is then added to this solution, and pure copper metal precipitates out because of the following chemical reaction: Fe(s) + CuSO4(aq) → Cu(s) + FeSO4(aq) Suppose an industrial quality-control chemist analyzes a sample from a copper processing plant in the following way. He adds powdered iron to a 150 mL copperIt) sulfate sample from the plant until no more copper will precipitate. He then washes, dries, and weighs the precipitate, and finds that it has a mas of 127. mg Caiculate the original concentration of coppert!I) sulfate in the sample. Round your answer to 3 significant digits
Answer #1

Atomic mass of Cu is 63.6 g/mol.

The mass of Cu precipitate obtained is 127. mg. Convert mg into g.

\dfrac { \text { 127. mg } }{ \text { 1000 mg/g } } = \text { 0.127 g }

The number of moles of Cu

\dfrac { \text { 0.127 g } }{ \text { 63.6 g/mol } } = \text { 0.001997 mol }

0.001997 moles of Cu are obtained from 0.001997 moles of copper sulphate.

[CuSO_4]=\dfrac { \text { 0.001997 mol }}{ \text { 150. mL} } \times \text { 1000 mL/L } = \text { 0.0133 mol/L }